A write 4.0 m long and 6.0 mm in a diameter has a resistance of 96 . If a potential difference of 20 V is applied between the ends, (a) what is the current in the wire? (b) what is the current density? (c) what is the resistivity of the wire material? 


1
Expert's answer
2021-11-01T19:24:18-0400

(a) By Ohm's law we get

"I=\frac{U}{R} \\

I = \frac{20}{96 \times 10^{-3}} = 208.33 \;A"

(b) If we suppose that the current density j is uniform, we get by definition jS =I , where S is a section area. As section is circular

"S = \pi r^2 = \pi (\frac{d}{2})^2\\

j = \frac{I}{S}\\

 = \frac{208.33}{3.14 \times 9 \times 10^{-6}} \\

= 7.37 \times 10^6 \;A/m^2"

(c) By definition of resistivity "R = \frac{\rho L}{S}"

"\rho" is material resistivity

L is wire length

"\rho = \frac{RS}{L} \\

= \frac{96 \times 10^{-3} \times 3.14 \times 9 \times 10^{-6}}{4.0} \\

= 6.78 \times 10^{-7} \; \Omega \cdot m"